### v1

# Empirical Stability Analysis of Kolmogorov-Arnold Networks in Hard-Constrained Recurrent Physics-Informed Discovery

Abstract
We investigate the integration of Kolmogorov-Arnold Networks (KANs) into hard-constrained recurrent physics-informed architectures (HRPINN) to evaluate the fidelity of learned residual manifolds in oscillatory systems. Motivated by the Kolmogorov-Arnold representation theorem and preliminary gray-box results, we hypothesized that KANs would enable efficient recovery of unknown terms compared to MLPs. Through initial sensitivity analysis on configuration sensitivity, parameter scale, and training paradigm, we found that while small KANs are competitive on univariate polynomial residuals (Duffing), they exhibit severe hyperparameter fragility, instability in deeper configurations, and consistent failure on multiplicative terms (Van der Pol), generally outperformed by standard MLPs. These empirical challenges highlight limitations of the additive inductive bias in the original KAN formulation for state coupling and provide preliminary empirical evidence of inductive bias limitations for future hybrid modeling.
1 Introduction
Hard-constrained recurrent architectures, such as the Hybrid Recurrent Physics-Informed Neural Network (HRPINN) (Spotorno et al., 2025), embed known physics structurally within a recurrent integrator, forcing the network to learn only residual dynamics. This design ensures physical consistency by construction and was demonstrated to be effective for accuracy and invariant enforcement in cyber-physical systems. A key open direction identified in that work is the symbolic discovery of unknown residual terms. Recently, Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024) have emerged as a promising alternative for scientific machine learning. Grounded in the Kolmogorov-Arnold representation theorem, which decomposes multivariate functions into sums of univariate functions ($\Phi(\mathbf{x})=\sum_{q}\phi_{q}(\sum_{p}\psi_{q,p}(x_{p}))$), the original KAN formulation replaces fixed activations in MLPs with learnable univariate B-splines. This additive structure provides a strong inductive bias for physical laws that are often expressed as truncated series expansions of non-linear terms (Liu et al., 2024). Notably, a shallow KAN can exactly represent any univariate polynomial, such as the cubic term in the Duffing oscillator, with sufficient spline resolution. In gray-box settings, KANs have shown potential in recovering hidden source terms in Neural ODEs (Koenig et al., 2024; Ma et al., 2025; Daryakenari et al., 2026).
Recent data-driven extensions enable symbolic residual recovery without priors, such as Structured Kolmogorov-Arnold Neural ODEs (SKANODEs) (Liu et al., 2025) for interpretable latents and symbolic equations from indirect data, and KAN-ODEs (Koenig et al., 2024) for hidden physics in ODEs/PDEs. Our work baselines the original vanilla KAN in recurrent physics-constrained settings like HRPINN. We hypothesized that replacing HRPINN’s MLP residual branch with a KAN would yield superior discovery accuracy and parameter efficiency, particularly because the additive bias should naturally isolate independent physical contributions. To test this hypothesis, we selected two canonical oscillators with contrasting residual structures: the Duffing oscillator, whose residual is a univariate polynomial ($0.3x^{3}$), and the Van der Pol oscillator, whose residual involves a multiplicative interaction ($(1-x^{2})v$). These oscillators were selected to represent the boundary between additive separability (Duffing) and multiplicative coupling (Van der Pol). Although KANs can theoretically represent multiplication through composition (for instance, $xy=\frac{1}{4}((x+y)^{2}-(x-y)^{2})$), this requires deeper layers, raising questions about practical stability in a recurrent setting. Through carefully controlled studies, we establish a baseline for the suitability of vanilla KANs in hard-constrained recurrent architectures. While the hypothesis holds for univariate terms, our experiments reveal substantial challenges for variable interaction.
2 Experimental Methods and Results
We utilize the HRPINN framework, in which the residual branch $R_{\theta}(x,v)$ receives the normalized state $[x,v]$ and is implemented either as a standard ReLU MLP or as a B-spline KAN. The known physical dynamics are hard-coded into the recurrent update rule. Training is performed either with single-step teacher forcing or with backpropagation through time (BPTT). Performance is evaluated on held-out trajectories using test MSE and Discovery $R^{2}$, which measures grid-based correlation with the true residual. Rather than using KAN-specific symbolic pruning, we utilize a unified candidate-based fitting approach for both KAN and MLP. This assesses the accuracy of surface recovery independently of the symbolic extraction mechanism, providing a direct comparison of inductive bias effectiveness. We conducted 3 complementary studies, always with 100 seeds, to probe different aspects of KAN behavior: a configuration ablation varying grid size and sparsity to assess hyperparameter stability, a parameter-scale ablation under teacher forcing to compare efficiency, and a parameter-scale ablation with BPTT to examine the effect of recurrent unrolling. Discovery $R^{2}$ is calculated on a dense $100\times 100$ grid over the phase space $x,v\in[-2.5,2.5]$, comparing the network’s predicted residual $R_{\theta}(x,v)$ against the analytical truth.
Table 1: Configuration Ablation (95% Bootstrap CI, $N=100$ seeds). $G$ and $k$ denote grid size and spline order. Config A ($G=5,k=3$) and Config F ($G=3,k=3$) represent the most stable KAN baselines. Many configurations fail severely on Van der Pol.
Configuration Duffing Discovery $R^{2}$ VdP Discovery $R^{2}$
Config A (Baseline) 0.835 $\pm$ 0.030 0.667 $\pm$ 0.037
Config B (Spline-Forced) 0.773 $\pm$ 0.097 0.320 $\pm$ 0.330
Config C (Sparse-Low) 0.595 $\pm$ 0.033 -5.229 $\pm$ 5.091
Config D (Sparse-High) 0.582 $\pm$ 0.026 -1.688 $\pm$ 1.318
Config E (Aggressive-Grid) 0.794 $\pm$ 0.067 0.699 $\pm$ 0.065
Config F (Coarse-Grid) 0.862 $\pm$ 0.037 0.639 $\pm$ 0.302
Config G (Fine-Grid) 0.745 $\pm$ 0.099 -0.174 $\pm$ 0.691
MLP (Small, 337 params) 0.957 $\pm$ 0.009 0.768 $\pm$ 0.015
We begin by examining the stability of KAN training across hyperparameter configurations. Table 1 reports Discovery $R^{2}$ for seven grid configurations over 100 random seeds. While certain coarse-grid settings achieve strong performance on the Duffing oscillator, surpassing the baseline in some cases, most configurations yield high variance or outright failure on the Van der Pol oscillator, with several producing negative $R^{2}$ values indicative of diverging solutions. In contrast, MLPs exhibited robust performance across standard settings. This extreme sensitivity underscores a practical fragility in the original KAN formulation that is absent in conventional MLPs (Noorizadegan et al., 2025), justifying our choice of oscillators as representative test cases for stability. Having identified viable configurations, we next compared parameter efficiency under teacher forcing. As shown in Table 2, very small KANs ($\approx$ 120 parameters) perform competitively with similarly sized MLPs on the univariate Duffing residual and occasionally outperform larger MLPs on this task, supporting our hypothesis for separable terms. However, the picture changes for Van der Pol: even modestly wider or deeper KANs collapse, yielding near-zero or negative $R^{2}$, while MLPs scale gracefully with parameter count and consistently achieve higher accuracy.
Switching to full recurrent training with BPTT partially mitigates these issues for shallow KANs (Table 2). The smallest KAN configuration achieves its highest Van der Pol score under BPTT ($R^{2}$ $\approx$ 0.74), suggesting that longer horizon supervision helps stabilize learning of variable interaction. Nevertheless, MLPs still dominate across nearly all scales, and deeper KANs remain catastrophically unstable, as evidenced by the high standard deviation and training failures. Qualitative examination of learned residual surfaces reinforces these quantitative findings (Figure 1). For the Duffing oscillator, the median small KAN accurately reproduces the characteristic cubic shape; simplified candidate fitting yielded $R^{2}=0.91$ for a discovered residual of $0.234x^{3}$ (Ground Truth: $0.3x^{3}$). While the coefficient is underestimated, the KAN captured the cubic structure in 38% of seeds, outperforming the MLP’s fitted form ($R^{2}=0.85$) in terms of local correlation. For Van der Pol, however, the same KAN approximates the multiplicative structure poorly, often collapsing to a roughly linear form rather than the expected parabolic modulation.
Table 2: Parameter ablation — One-Step Teacher Forcing vs BPTT (Mean $\pm$ 95% CI, N=100 seeds).
Arch. Config. Params One-Step TF BPTT
Duffing $R^{2}$ VdP $R^{2}$ Duffing $R^{2}$ VdP $R^{2}$
KAN Very Small 120 0.836 $\pm$ 0.032 0.464 $\pm$ 0.166 0.914 $\pm$ 0.061 0.743 $\pm$ 0.061
Small 240 0.777 $\pm$ 0.079 0.322 $\pm$ 0.292 0.874 $\pm$ 0.080 0.785 $\pm$ 0.073
Wide 480 0.845 $\pm$ 0.025 0.232 $\pm$ 0.570 0.468 $\pm$ 0.773 -0.602 $\pm$ 2.842
Deep 880 -3.146 $\pm$ 7.106 -0.303 $\pm$ 1.579 (Unstable) 0.754 $\pm$ 0.079
MLP Tiny 105 0.914 $\pm$ 0.026 0.593 $\pm$ 0.048 0.906 $\pm$ 0.092 0.622 $\pm$ 0.173
Small 337 0.957 $\pm$ 0.009 0.768 $\pm$ 0.015 0.937 $\pm$ 0.047 0.879 $\pm$ 0.032
Medium 1185 0.960 $\pm$ 0.013 0.805 $\pm$ 0.014 0.951 $\pm$ 0.033 0.879 $\pm$ 0.019
Large 4417 0.965 $\pm$ 0.009 0.843 $\pm$ 0.010 0.932 $\pm$ 0.063 0.898 $\pm$ 0.017
Note: Mean $\pm$ 95% CI, N=100 seeds. “(Unstable)” indicates training instability observed for deeper KANs under BPTT.
Figure 1: Comparison of True Residual Surfaces vs. KAN-Recovered Manifolds (Config A). (a) Duffing: The KAN successfully identifies the univariate cubic manifold ($x^{3}$), demonstrating high fidelity in separable terms. (b) Van der Pol: The KAN struggles to resolve the multiplicative structure ($(1-x^{2})v$), approximating the magnitude (with relevant error given the difference in the colorbars) but failing on the coupled interaction. Results show representative seeds (65 and 82) that align with the median performance reported in Table 1.
3 Discussion and Conclusion
Our experiments offer a nuanced perspective on the original KAN formulation within hard-constrained physics-informed recurrent architectures. The hypothesis is validated for separable, univariate residuals: small KANs efficiently recover polynomial non-linearities with parameter counts competitive to MLPs. Yet a critical limitation emerges for multiplicative interactions. Although KANs can theoretically express multiplication through layered composition, our deeper vanilla configurations (Table 2, ”Deep” row) prove highly unstable during recurrent optimization (Sohail, 2024), a problem exacerbated by rapid error accumulation in the hard-constrained integration loop. MLPs, utilizing dense matrix multiplication, enforce variable interaction in the first layer ($w_{i}x+w_{j}v$). In contrast, vanilla KANs possess a strong additive inductive bias ($\phi(x)+\phi(v)$). To model the multiplicative Van der Pol term $(1-x^{2})v$, the KAN must learn to approximate the product operator through deep composition. Our results demonstrate that even with the correct functional form provided via candidate fitting, the KAN fails to accurately recover the multiplicative manifold. This suggests that the primary bottleneck in recurrent KAN integration is the optimization stability of the composition, not the symbolic extraction process itself, leading to the observed optimization challenges in variable interaction.
We note that subsequent KAN variants, such as deeper operator networks, hybrid formulations, or specialized scaling for oscillatory dynamics (Abueidda et al., 2025; Zhang et al., 2025; Mostajeran and Faroughi, 2025), may overcome these shortcomings. For instance, SKANODEs (Liu et al., 2025) and KAN-ODEs (Koenig et al., 2024) demonstrate symbolic potential in continuous Neural ODEs, motivating extensions to recurrent HRPINN for state coupling. The present study utilizes a unified candidate-fitting benchmark to assess manifold fidelity. While this levels the comparison with MLPs, it is important to note that KANs possess a unique structural advantage: the potential for direct symbolic recovery through spline pruning, a capability entirely absent in standard MLPs. Our results suggest that while this latent advantage exists, its realization in recurrent physics-informed settings is currently throttled by optimization instability rather than a lack of expressivity. While deeper or hybrid KAN variants may improve expressivity for multiplicative interactions, they typically incur higher computational costs and can introduce additional training instability, which we leave for future work. Therefore, future investigations should target stabilized learning of interaction terms and integrate automatic symbolic extraction to fully realize the interpretive potential of KAN-based physics-informed modeling, building on this initial PoC baseline for vanilla KANs in recurrent settings.
Acknowledgments
This work was partially supported by FUNDEP grants Rota 2030/Linha VI 29271.02.01/2022.01-00 and 29271.03.01/2023.04-00.